# ACT-R: Adaptive Camera Trajectories for 3D Reconstruction from Single Image

### v1

@@ Abstract
We introduce adaptive view planning to multi-view synthesis, aiming to improve both occlusion revelation and 3D consistency for single-view 3D reconstruction. Instead of generating an unordered set of views independently or simultaneously, we generate a sequence of views, leveraging temporal consistency to enhance 3D coherence. Most importantly, our view sequence is not determined by a pre-determined camera setup. Instead, we compute an adaptive camera trajectory (ACT), specifically, an orbit of camera views, which maximizes the visibility of occluded regions of the 3D object to be reconstructed. Once the best orbit is found, we feed it to a video diffusion model to generate novel views around the orbit, which in turn, are passed to a multi-view 3D reconstruction model to obtain the final reconstruction. Our multi-view synthesis pipeline is quite efficient since it involves no run-time training/optimization, only forward inferences by applying the pre-trained models for occlusion analysis and multi-view synthesis. Our method predicts camera trajectories that reveal occlusions effectively and produce consistent novel views, significantly improving 3D reconstruction over SOTA on the unseen GSO dataset, both quantitatively and qualitatively.
Figure 1: Our single-view 3D reconstruction method, ACT-R, predicts an adaptive camera trajectory (green) to maximize the visibility of occluded object parts within a fixed sequence length, achieving this in under 10s. The trajectory is then used by a video generator (e.g., SV3D [60]) to produce a sequence of novel views for multi-view 3D reconstruction, here using NeuS [22]. Compared to a generic trajectory (red, at fixed elevation), ACT-R yields much cleaner results with more faithful recovery of occluded regions.
@@ 1 Introduction
Figure 2: Pipeline of our single-view reconstruction method, ACT-R, with adaptive camera trajectories (ACT). We first employ Slice3D [64] to produce the slice images of the input object, with the slicing direction from the camera to the object center. Then we compute the semantic difference between the input and its slices by comparing their $512\times 7\times 7$ feature maps extracted from VGG16 [51]. For better visualization, each difference map $d_{i}\in[0,1]^{7\times 7}$ is up-scaled and overlaid onto slice images . Afterwards, we identify the regions that have significant semantic differences (Sec. 3.2), and plan the camera trajectories based on them (Sec. 3.3). Finally, we condition SV3D [60] on our planned trajectories, yielding a sequence of views, which can be fed into NeuS [62] and LRM [71] for 3D reconstruction.
@@
3D reconstruction from single-view images has been one of the most intensively studied problems in computer vision. One class of modern approaches directly generate or regress 3D representations of objects [75, 30, 69, 19, 58] from input images. With significant advances in both novel view synthesis (e.g., neural radiance field or NeRF [42] and 3D Gaussian splatting [24]) and multi-view 3D reconstruction (e.g., NeuS [62] and variants [36, 61]), the other line of popular approaches to single-view 3D reconstruction first perform a multi-view synthesis, which typically generates an unordered set of views either independently [33] or simultaneously [35, 37, 56, 50] with fixed camera setups.
@@
The main challenges to single-to-multi view synthesis are twofold. First, the multi-view images should reveal the occluded structures of the target 3D object that are hidden from the input image, which is a highly ill-posed problem. Second, the generated views must attain 3D consistency to ensure that a plausible and coherent 3D model can be reconstructed. Upon close examination of state-of-the-art multi-view synthesis methods, as well as single-view reconstruction methods which combine such syntheses with direct 3D prediction [26, 66, 54, 32], we find that there is still much room for improvement in addressing the two challenges.
@@
In this paper, we introduce adaptive view planning to multi-view synthesis, so as to improve both occlusion revelation and 3D consistency for single-view 3D reconstruction. Instead of synthesizing an unordered set of views as in prior works, we generate a sequence of views, leveraging the inherent temporal consistency to enhance 3D coherence. Most importantly, our view sequence is not determined by a pre-determined [35, 37, 56, 41, 72]. Instead, we compute an adaptive camera trajectory (ACT), which is specific to the target 3D object and its input view, to generate the sequence of novel views. Note that several recent works have shown benefits of camera perturbation [60] and stochastic conditioning [67] for novel view synthesis. Our approach takes these further with judicious view planning.
@@
Once the best orbit is found, we feed it to a video diffusion model to generate a sequence of novel views that are adapting to the 3D object. The multi-view images obtained are finally passed to a 3D reconstruction model to obtain the final result. These last two steps can employ any state-of-the-art video diffusion and multi-view 3D reconstruction models, while for our current work, we employ Stable Video 3D (SV3D) [60] and NeUS [62], respectively.
Fig. 2 shows the entire pipeline of our method, which is coined ACT-R for using Adaptive Camera Trajectory for single-view 3D Reconstruction. Our multi-view generation pipeline, prior to NeuS or LRM, is quite efficient since it involves no run-time training/optimization, only forward inferences by applying the pre-trained Slice3D and SV3D.
Extensive experiments demonstrate that our method predicts camera trajectories tailored to each input example, effectively revealing occluded regions for high-quality reconstruction. On the GSO benchmark [9], our approach outperforms state-of-the-art methods, including direct image-to-3D [27] and those based on multi-view syntheses from static [60], fixed [53, 37], and randomly perturbed [60] camera trajectories, both qualitatively and quantitatively.
@@ 2.1 Video Generation in 2D and for 3D
Early works [17, 52, 16] extend image diffusion to video generation. Stable Video Diffusion (SVD) [3] adapts latent diffusion methods [48, 4] to large-scale video datasets for temporally coherent generation. A stream of work focuses on keyframe conditioning [12, 63, 28, 74], where initial frames are generated to anchor subsequent video synthesis, with latent-consistency networks ensuring temporal and appearance coherence. PYoCo [11] employs tailored noise priors for fine-tuning image models, while Lumiere [2] represents videos in a single latent space without keyframe dependencies. Methods like Tune-a-video [68] enable few-shot video synthesis through fine-tuning of pre-trained image diffusion models. Training-free approaches [18, 21] utilize LLMs for generation guidance, while additional works [77, 78] have explored various latent diffusion strategies for video synthesis. Although video generation models typically lack explicit 3D representations, they can still achieve 3D consistency by producing temporally coherent videos. In our approach, we use video diffusion models as the backbone to generate sequences of views.
@@ 2.2 Multi-view Synthesis from Single View
Generating multiple views from a single image in one step presents significant consistency challenges, as establishing correspondences between substantially different viewpoints remains difficult. To address this issue, several approaches have been proposed. 3DiM [67] generates novel views in an auto-regressive manner, selecting a previously generated view as a condition for producing each subsequent view during the denoising process. ViewFusion [72] builds upon Zero-1-to-3 [33] to generate novel views using a similar auto-regressive approach as 3DiM [67]. IM-3D [41] employs a video generation model [12] to create novel views that are then processed by 3D G-Splat [24] for 3D reconstruction. This approach can be iteratively refined by feeding rendered objects back into the video diffusion model. However, these methods (3DiM, ViewFusion, IM-3D) rely on either random camera poses or predefined trajectories for novel view generation, without considering the specific structural properties of the objects being modeled.
@@ 2.3 View and Path Planning
Path planning is an important task with applications in navigation, scanning, and even computational fabrication [10, 23, 46, 47, 34, 38]. In particular, Next-Best-View (NBV) planning addresses the fundamental challenge of determining an optimal sequence of camera positions to maximize information gain during scene or object inspection  [8, 40, 59]. Two popular traditional approaches to this problem include voxel-space methods that optimize coverage metrics  [8, 40, 59, 39], and surface-based methods that analyze boundary characteristics to determine optimal viewpoints  [44, 25, 6]. Recent advances in deep learning have transformed the NBV paradigm. Daryl et al. [43] leveraged reinforcement learning to predict optimal viewpoints by maximizing coverage-based rewards. Liu et al. [34] developed a learning-based reconstructability predictor to guide their iterative view planner for 3D reconstruction. NeU-NBV [22] introduced a novel uncertainty evaluation framework that utilizes neural image rendering to inform NBV prediction. GenNBV [7] proposed a generalizable NBV policy to enhance its performance on unseen scenes.
While traditional NBV approaches assume the availability of a complete 3D reference model or scene, our work addresses a more challenging scenario where only a single image serves as input. This constraint fundamentally shifts the problem from pure coverage optimization to view prediction based on limited initial information, requiring novel strategies for viewpoint selection and trajectory planning.
@@ 3 Method
Similar to SV3D [60], we assume that the camera always looks at the center of an object (origin of the world), and the distance between the camera and the center remains unchanged, so any viewpoint can be specified by only two parameters: elevation and azimuth, $\pi_{i}=(e_{i},a_{i})$, where $e_{i},a_{i}$ are the elevation and azimuth angles.
When designing an adaptive trajectory in contrast with a generic one, we aim to avoid increasing its length (i.e., the number of views), as doing so would complicate both the multi-view generation and the subsequent 3D reconstruction processes. To maintain consistency, we fix the number of generated views, $N$, as $21$ and use a single closed orbit for the camera trajectory, the same as SV3D (u) [60] for a fair comparison as SV3D serves as our closest baseline.
@@ 3.1 Overview
From these blocks, we plan a camera trajectory aiming to maximize the visibility of the occluded regions, with each block weighted by its semantic difference. This optimized trajectory is fed into SV3D [60] to generate a sequence of novel views. The generated novel views can be used in a plug-and-play manner for any 3D reconstruction pipeline that takes novel views as intermediate representations. We used Neus [62] and large reconstruction models [71] as our two alternative 3D reconstruction backbones.
@@ 3.2 Building Semantic Difference Blocks
Figure 3: Left: Transforming the semantic difference maps into 3D blocks, where lighter yellow indicates greater differences. Right: Visualization of different camera orbits. Red: fixed elevation; Green: variable elevations that capture greater visibility. “diff” and “cam” denote difference and camera, respectively.
As illustrated in the first step in Fig. 2, Slice3D produces in total $M$ slice images $\{s_{1},s_{2},\dots,s_{M}\}$ from the input $x$. We quantify the semantic differences between slice image $s_{i}$ and input $x$ by comparing the features from the final pooling layer of VGG16 [51]. Specifically, we compute:
$d_{i}=\langle\phi(x),\phi(s_{i})\rangle,$ (1)
@@ 3.3 Camera Trajectory Planning
The camera is initially positioned at the input view, i.e., $(e_{0},a_{0})=(0,\alpha)$. The generic camera trajectory from SV3D (u) [60] uses a fixed elevation $\alpha$, shown as the red orbit in Fig. 3. $r$ and $\alpha$ are used to calculate the world-to-camera coordinate transform since SV3D [60] operates in a world frame while we use a camera coordinate system.
The reconstructed coordinate system is shown in Fig. 3 (right). To maximize the observation of occluded regions, an ideal trajectory should aim to cover as many blocks as possible, prioritizing those blocks with greater semantic differences (denoted by lighter-coloured blocks). With the camera’s known field of view (FOV) $\theta$, we can determine the block visibilities from any given position.
@@
where $\psi(\pi(t))$ denotes the set of all visible blocks under camera $\pi(t)$. More details about the $\psi(\cdot)$ formulation can be found in the supplementary material.
@@ 4 Experiments
Evaluation Metrics. For 3D metrics, we use F-score%1 (F1) [57], Chamfer $\mathcal{L}_{2}$ (CD), Hausdorff (HD), and Light Field distances (LFD) [5] to evaluate reconstruction results. In addition, we also report 2D metrics such as PSNR, SSIM [65], and LPIPS [76] on 12 rendered views of the reconstructed meshes. Since no single metric is entirely informative, we provide a comprehensive evaluation to cover both object- and image-space, as well as distortion and visual similarity aspects of the 3D reconstruction.
@@ 4.1 Qualitative Visual Comparisons
We compare ACT-R to representative SOTA approaches for single-view 3D reconstruction, including Wonder3D [37], SV3D(u) [60] for volumetric-rendering-based reconstruction, LGM [53] based on LRMs, and CraftsMan [27] as one of the latest direct image-to-3D methods. We further present three ablated experiment setups:
@@
    $\text{Ours}_{\text{lrm}}$: LRM with 6 views sampled uniformly from adaptive camera trajectory produced by our method.
Qualitative results in Fig. 4 show that different reconstruction backbones offer distinct advantages, while also expose their limitations. NeuS produces more faithful reconstructions that closely match the input, yet suffers from blobby surface artifacts. Effectively leveraging 3D priors, LRMs tend to produce smoother surfaces, but respect less the details from the input image. Despite these inherent backbone-specific characteristics, ACT-R’s adaptive trajectories consistently improve reconstruction quality, particularly over occluded regions. Our approach more faithfully reconstructs the solid bowl interiors (rows 2 and 3) that most competing methods overlooked due to occlusion in the input image. It also provides enhanced spatial awareness between objects (row 5) and creates reasonable joint connections in complex structures such as the umbrellas (row 6).
Fig. 5 compares the generated views of Wonder3D [37] and SV3D [60] to ours. Compared to SV3D, Wonder3D offers a wider range of visibility (e.g., bottom views of the objects) by varying both camera azimuths and elevations. However, it tends to suffer from multi-view inconsistency, as seen in the heavily distorted back view. SV3D maintains relatively good consistency but cannot observe the back of the object, which will lead to inadequate information during 3D reconstruction. In contrast, our method captures the bottom of the object and achieves view consistency.

### v2

# ACT-R: Adaptive Camera Trajectories for Single-View 3D Reconstruction

@@ Abstract
We introduce the simple idea of adaptive view planning to multi-view synthesis, aiming to improve both occlusion revelation and 3D consistency for single-view 3D reconstruction. Instead of producing an unordered set of views independently or simultaneously, we generate a sequence of views, leveraging temporal consistency to enhance 3D coherence. More importantly, our view sequence is not determined by a pre-determined and fixed camera setup. Instead, we compute an adaptive camera trajectory (ACT), forming an orbit, which seeks to maximize the visibility of occluded regions of the 3D object to be reconstructed. Once the best orbit is found, we feed it to a video diffusion model to generate novel views around the orbit, which can then be passed to any multi-view 3D reconstruction model to obtain the final result. Our multi-view synthesis pipeline is quite efficient since it involves no run-time training/optimization, only forward inferences by applying pre-trained models for occlusion analysis and multi-view synthesis. Our method predicts camera trajectories that reveal occlusions effectively and produce consistent novel views, significantly improving 3D reconstruction over SOTA alternatives on the unseen GSO dataset.
Figure 1: Our single-view 3D reconstruction method, ACT-R, predicts an adaptive camera trajectory (green), aiming to maximize the visibility of occluded object parts over a fixed sequence length (20 views in this example). The trajectory, obtained in under 10s, is then used by a video generator (e.g., SV3D [1]) to produce a sequence of novel views for multi-view 3D reconstruction, here using the established NeUS model [2]. Compared to a generic trajectory (red), at fixed elevation, ACT-R yields much cleaner results with more faithful recovery of occluded regions.
@@ 1 Introduction
Figure 2: Pipeline of our single-view reconstruction method, ACT-R, with adaptive camera trajectories (ACT). We first employ Slice3D [3] to produce the slice images of the input object, with the slicing direction from the camera to the object center. Then we compute the semantic difference between the input and its slices by comparing their $512\times 7\times 7$ feature maps extracted from VGG16 [4]. For better visualization, each difference map $d_{i}\in[0,1]^{7\times 7}$ is up-scaled and overlaid onto slice images . Afterwards, we identify the regions that have significant semantic differences (Sec. 3.2), and plan the camera trajectories based on them (Sec. 3.3). Finally, we condition SV3D [1] on our planned trajectories, yielding a sequence of views, which can be fed into either NeUS [2] or InstantMesh (IM) [5] for multi-view 3D reconstruction.
@@
3D reconstruction from single-view images has been one of the most intensively studied problems in computer graphics. One class of modern approaches directly generate or regress 3D representations of objects from input images [6, 7, 8, 9, 10, 11], often resorting to 3D supervision which requires large 3D datasets for training. With significant advances in novel view synthesis [12, 13], the second line of popular approaches to single-view 3D reconstruction first perform a multi-view synthesis, which typically generates an unordered set of views either independently [14] or simultaneously [15, 16, 17, 18] with fixed camera setups. This is followed by multi-view 3D reconstruction via differentiable rendering (e.g., NeUS [2] and variants [19, 20]) so that the entire solution pipeline can avoid direct 3D supervisionWe use the term “direct 3D supervision” in the context of 3D reconstruction to refer to models trained using paired images and ground-truth 3D models..
@@
The main challenges to single-to-multi view synthesis are twofold. First, the multi-view images should reveal the occluded structures of the target 3D object that are hidden from the input image, which is a highly ill-posed problem. Second, the generated views must attain 3D consistency to ensure that a plausible and coherent 3D model can be reconstructed. Upon close examination of state-of-the-art multi-view synthesis methods, as well as single-view reconstruction methods which combine such syntheses with direct 3D prediction [21, 22, 23, 24], we find that there is still much room for improvement.
@@
In this paper, we introduce adaptive view planning to multi-view synthesis, so as to improve both occlusion revelation and 3D consistency for single-view 3D reconstruction. Instead of synthesizing an unordered set of views as in prior works, we generate a sequence of views, leveraging the inherent temporal consistency to enhance 3D coherence. More importantly, our view sequence is not constructed by a pre-determined setup [15, 16, 17, 25, 26]. Instead, we compute an adaptive camera trajectory (ACT), which is specific to the target 3D object and its input view, to generate the sequence of novel views. Note that several recent works have shown benefits of camera perturbation [1] and stochastic conditioning [27] for novel view synthesis. Our approach takes these further with judicious view planning.
@@
Once the best orbit is found, we feed it to a video diffusion model to generate a sequence of novel views that are adapting to the 3D object. The multi-view images obtained are finally passed to a 3D reconstruction model to obtain the final result. Note that these last two steps can employ a variety of state-of-the-art video diffusion and multi-view 3D reconstruction models. In our current work, we employ Stable Video 3D (SV3D) [1] for the former, while for the latter, either NeUS [2], a well established method, or InstantMesh (IM) [5], a more recent one based on large reconstruction models (LRMs). As a result, our entire solution pipeline (see Fig. 2), which is coined ACT-R for using Adaptive Camera Trajectory for single-view 3D Reconstruction, does not use direct 3D supervision since none of Slice3D, NeUS, or InstantMesh does. Also, our multi-view generation is quite efficient since it involves no run-time training or optimization, only forward inferences by applying the pre-trained Slice3D and SV3D.
Extensive experiments demonstrate that our method predicts camera trajectories tailored to each input example, effectively revealing occluded regions for higher-quality reconstruction. On the GSO benchmark [28], ACT-R outperforms, both qualitatively and quantitatively, state-of-the-art methods based on multi-view syntheses from static [1], fixed [23, 16], or randomly perturbed [1] camera trajectories, all without direct 3D supervision, as well as a recent method, Craftsman [29], which employs 3D supervision in the multi-view reconstruction step.
Figure 3: Initial single-view inputs to advanced methods such as Trellis and Hunyuan3D (HY) can be improved by one of the synthesized multi-view images from ACT-R, to improve reconstruction results; see highlights.
On the other hand, ACT-R, which is best positioned to boost multi-view 3D reconstruction methods without direct 3D supervision, cannot beat some of the latest LRMs such as CLAY [6], Hunyuan3D [10], or Trellis [9], which all use direct 3D supervision trained by closed-source (CLAY) or carefully curated (Trellis) large-scale 3D assets. Indeed, directly comparing ACT-R with these models can be unfair due to the different training setups. Nevertheless, as shown in Fig. 3, our ACT can benefit these supervised LRMs by improving the input views.
@@ 2 Related Work
As 3D reconstruction, even only for single views, has been extensively studied, we will only focus on closely related works to ACT-R, on video generation, multi-view synthesis, and view planning.
@@ 2.1 Video Generation in 2D and for 3D
Early works [30, 31, 32] extend image diffusion to video generation. Stable Video Diffusion (SVD) [33] adapts latent diffusion methods [34, 35] to large-scale video datasets for temporally coherent generation. A stream of work focuses on keyframe conditioning [36, 37, 38, 39], where initial frames are generated to anchor subsequent video synthesis, with latent-consistency networks ensuring temporal and appearance coherence. PYoCo [40] employs tailored noise priors for fine-tuning image models, while Lumiere [41] represents videos in a single latent space without keyframe dependencies. Methods like Tune-a-video [42] enable few-shot video synthesis through fine-tuning of pre-trained image diffusion models. Training-free approaches [43, 44] utilize Large Lanague Models (LLMs) for generation guidance, while other works [45, 46] have explored various latent diffusion strategies for video synthesis. Although video generation models typically lack explicit 3D representations, they can still achieve 3D consistency by producing temporally coherent videos. In our approach, we use video diffusion models as the backbone to generate view sequences.
@@ 2.2 Multi-view Synthesis from Single View
Generating multiple views from a single image in one step presents significant consistency challenges, as establishing correspondences between substantially different viewpoints remains difficult. To address this issue, 3DiM [27] generates novel views in an auto-regressive manner, selecting a previously generated view as a condition for producing each subsequent view during the denoising process. ViewFusion [26] builds upon Zero-1-to-3 [14] to generate novel views using a similar auto-regressive approach as 3DiM [27]. IM-3D [25] employs a video generation model [36] to create novel views that are then processed by 3D G-Splat [13] for 3D reconstruction. This approach can be iteratively refined by feeding rendered objects back into the video diffusion model. However, these methods (3DiM, ViewFusion, IM-3D) rely on either random camera poses or predefined trajectories for novel view generation, without considering the specific structural properties of the objects being modeled.
@@ 2.3 View and Path Planning
Path planning has applications in navigation, scanning, and even computational fabrication [54, 55, 56, 57, 58, 59]. In particular, Next-Best-View (NBV) planning addresses the fundamental challenge of determining an optimal sequence of camera positions to maximize information gain during scene or object inspection  [60, 61, 62]. Two popular traditional approaches to this problem include voxel-space methods that optimize coverage metrics  [60, 61, 62, 63], and surface-based methods that analyze boundary characteristics to determine optimal viewpoints  [64, 65, 66]. Recent advances in deep learning have transformed the NBV paradigm. Daryl et al. [67] leveraged reinforcement learning to predict optimal viewpoints by maximizing coverage-based rewards. Liu et al. [58] developed a learning-based reconstructability predictor to guide their iterative view planner for 3D reconstruction. NeU-NBV [68] introduced a novel uncertainty evaluation framework that utilizes neural image rendering to inform NBV prediction. GenNBV [69] proposed a generalizable NBV policy to enhance its performance on unseen scenes.
While traditional NBVs assume the availability of a complete 3D reference model, our work addresses a more challenging scenario where only a single image serves as input. This constraint fundamentally shifts the problem from pure coverage optimization to view prediction based on limited initial information, requiring novel strategies for viewpoint selection and trajectory planning.
@@ 3 Method
Similar to SV3D [1], we assume that the camera always looks at the center of an object (origin of the world), and the distance between the camera and the center remains unchanged, so any viewpoint can be specified by only two parameters: $\pi_{i}=(e_{i},a_{i})$, where $e_{i},a_{i}$ are the elevation and azimuth angles.
When designing an adaptive trajectory in contrast with a generic one, we aim to avoid increasing its length (i.e., the number of views), as doing so would complicate both the multi-view generation and subsequent 3D reconstruction. To maintain consistency, we fix the number of generated views, $N$, as $21$ and use a single closed orbit for the camera trajectory, the same as SV3D (u) [1] for a fair comparison as SV3D serves as our closest baseline.
@@ 3.1 Overview
From these blocks, we plan a camera trajectory aiming to maximize the visibility of the occluded regions, with each block weighted by its semantic difference. This optimized trajectory is fed into SV3D [1] to generate a sequence of novel views. The generated novel views can be used in a plug-and-play manner for any 3D reconstruction pipeline that takes novel views as intermediate representations. We used NeUS [2] and large reconstruction models trained from InstantMesh [5] as our two alternative 3D reconstruction backbones.
@@ 3.2 Building Semantic Difference Blocks
Figure 4: Illustration of camera trajectory planning. Left: Transforming the semantic difference maps into 3D blocks, where lighter yellow indicates greater differences. Right: Visualization of different camera orbits. Red: fixed elevation; Green: variable elevations that capture greater visibility. “diff” and “cam” denote difference and camera, respectively.
@@
Compared with more recent feature encoders such as DINOv2 [70], we empirically find that VGG16 [4] works better in extracting semantic difference maps between slices and input images. Please see the supplementary materials for detailed ablation studies.
@@ 3.3 Camera Trajectory Planning
The camera is initially positioned at the input view, i.e., $(a_{0},e_{0})=(0,\alpha)$. The generic camera trajectory from SV3D (u) [1] uses a fixed elevation $\alpha$, shown as the red orbit in Fig. 4 (right). Here $r$ and $\alpha$ are used to calculate the world-to-camera coordinate transform since SV3D [1] operates in a world frame while we use a camera coordinate system.
The reconstructed coordinate system is shown in Fig. 4. To maximize the observation of occluded regions, an ideal trajectory should aim to cover as many blocks as possible, prioritizing those blocks with greater semantic differences (denoted by lighter-coloured blocks). With the camera’s known field of view (FOV) $\theta$, we can determine the block visibilities from any given position.
@@
where $\psi(\pi(t))$ denotes the set of all visible blocks under camera $\pi(t)$. More details about the $\psi(\cdot)$ formulation and orbital camera trajectory justification can be found in the supplementary material.
@@ 4 Experiments
Evaluation Metrics. For 3D metrics, we use Chamfer $\mathcal{L}_{2}$ (CD), Hausdorff (HD), and Light Field distances (LFD) [74], as well as F-score%1 (F1) [75], to evaluate reconstruction results. In addition, we also report 2D metrics such as PSNR, SSIM [76], and LPIPS [77] on 12 rendered views of the reconstructed meshes. Since no single metric is entirely informative, we provide a comprehensive evaluation to cover both object- and image-space, as well as distortion and visual similarity aspects of the 3D reconstruction.
@@ 4.1 Qualitative Visual Comparisons
We compare ACT-R to representative SOTA approaches for single-view 3D reconstruction, including Wonder3D [16], SV3D(u) [1] for volumetric-rendering-based reconstruction, LGM [23] based on LRMs, and CraftsMan [29] as an image-to-3D method with direct 3D supervision. We further present three ablated experiment setups:
@@
    $\text{Ours}_{\text{IM}}$: LRM with 6 views sampled uniformly from adaptive camera trajectory produced by our method.
Qualitative results in  Figures 9 and 10 show that different reconstruction backbones offer distinct advantages, while also expose their limitations. NeUS produces more faithful reconstructions that closely match the input, yet suffers from blobby surface artifacts. Effectively leveraging 3D priors, LRMs produce smoother surfaces, but respect less the details from the input image. Despite these inherent backbone-specific characteristics, ACT-R’s adaptive trajectories consistently improve reconstruction quality, particularly over occluded regions. Our approach more faithfully reconstructs the solid bowl interiors (rows 2 and 3) that most competing methods overlooked due to occlusion in the input image. It also provides enhanced spatial awareness between objects (row 5) and creates reasonable joint connections in complex structures such as the umbrellas (row 6).
Fig. 7 compares the generated views of Wonder3D [16] and SV3D [1] to ours. Compared to SV3D, Wonder3D offers a wider range of visibility (e.g., bottom views of the objects) by varying both camera azimuths and elevations. However, it tends to suffer from multi-view inconsistency, as seen in the heavily distorted back view. SV3D maintains better consistency but cannot observe the back of the object, leading to inadequate information during 3D reconstruction. In contrast, our method captures the bottom of the object and achieves view consistency.